### v1

# See More Details: Efficient Image Super-Resolution by Experts Mining

Abstract
Reconstructing high-resolution (HR) images from low-resolution (LR) inputs poses a significant challenge in image super-resolution (SR). While recent approaches have demonstrated the efficacy of intricate operations customized for various objectives, the straightforward stacking of these disparate operations can result in a substantial computational burden, hampering their practical utility. In response, we introduce SeemoRe, an efficient SR model employing expert mining. Our approach strategically incorporates experts at different levels, adopting a collaborative methodology. At the macro scale, our experts address rank-wise and spatial-wise informative features, providing a holistic understanding. Subsequently, the model delves into the subtleties of rank choice by leveraging a mixture of low-rank experts. By tapping into experts specialized in distinct key factors crucial for accurate SR, our model excels in uncovering intricate intra-feature details. This collaborative approach is reminiscent of the concept of “see more”, allowing our model to achieve an optimal performance with minimal computational costs in efficient settings. The source will be publicly made available at https://github.com/eduardzamfir/seemoredetails
Mixture of Experts, Efficiency, Image Super-Resolution
1 Introduction
002020202040404040606060608080808010010010010012012012012014014014014016016016016018018018018020020020020022022022022024024024024026026026026028028028028038.738.738.738.738.838.838.838.838.938.938.938.93939393939.139.139.139.139.239.239.239.239.339.339.339.339.439.439.439.439.539.539.539.5+$0.18$dB-$51\%$+$0.16$dB-$17\%$ better CNN Transformers ShuffleMixerIMDNPANDDistill-SRSAFMNSeemoRe-TSeemoRe-BSeemoRe-LSRFormerSwinIRGMACSPSNR (dB)×$\times 2$ SR on Manga109
Figure 1: Model complexity trade-off. Visualization of PSNR, GMACS, and parameter counts on Manga109 dataset for $\times$2 task. Our proposed SeemoRe excels the state-of-the-art CNN-based and lightweight Transformer-based SR models. Marker size indicates parameter counts w.r.t SwinIR-Light (Liu et al., 2021).
Single image super-resolution (SR) is a long-standing low-level vision endeavour that pursues the reconstruction of a high-resolution (HR) image from its degraded low-resolution (LR) counterpart. This challenging task has garnered considerable attention owing to the expeditious development of ultra-high definition devices and video streaming applications  (Khani et al., 2021; Zhang et al., 2021a). Foreseeing the resource constraints, it is of substantial desire to design an efficient SR model for gauging the HR images to be perfectly visualized on these devices or platforms. Identifying the most plausible candidates for missing HR pixels poses a particular challenge for SR. In the absence of external priors, the primary approaches for SR involves exploring the intricate relationships among the neighboring pixels for reconstruction. Recent SR models exemplify this through methods such as (a) attention  (Liang et al., 2021; Zhou et al., 2023; Chen et al., 2023), (b) feature mixing  (Hou et al., 2022; Sun et al., 2023), and (c) global-local context modeling  (Wang et al., 2023; Sun et al., 2022), yielding remarkable accuracy.
Unlike other approaches in this work, we aim to avoid complex and disconnected blocks focusing on specific factors, opting instead for a unified learning module specialized for all aspects. However, an additional challenge arises due to the efficiency requirement, rendering implicit learning through a vast number of parameters unfeasible, especially in the context of devices with limited resources.
To achieve such an efficient unification, we introduce SeemoRe, which leverages the synergy of different experts to maximize intra-feature intertwining, collaboratively learning a cohesive relation across LR pixels. Our motivation stems from the observation that image features often display diverse patterns and structures. Attempting to capture and model all these patterns with a single, monolithic model can be challenging. Collaborative experts, on the other hand, enable the network to specialize in different regions or aspects of the input space, enhancing its adaptability to various patterns and facilitating the modeling of LR-HR dependencies, akin to “See More”.
Technically, our network is composed of stacked residual groups (RGs) for dynamically selecting the pivotal features via experts, focusing on two different aspects. At the macro level, each RG embodies two successive expert blocks: (a) Rank modulating expert (RME), expertized in dealing with the most informative features through low-rank modulation, and (b) Spatial modulating expert (SME), expertized in efficient spatial enhancement. At the micro level, we devise a Mixture of Low-Rank Expertise (MoRE) as the foundational component within RME to dynamically select the best and most suitable rank for different inputs and at different network depths while implicitly modeling the global contextual relationships. Furthermore, we design a Spatial Enhancement Expertise (SEE) as an efficient alternative to complex self-attention within SME for distinctly improving the spatial-wise local aggregation capabilities. Such a combination efficiently modulates the mutual dependencies within the feature attributes, enabling our model to extract high-level information, which is a key aspect of SR. By explicitly mining experts at different granularity for different expertise, our network navigates the intricacies between spatial and channel features, maximizing their synergistic contribution and thus accurately and efficiently reconstructing more details.
As shown in Figure 1, our network significantly outperforms the state-of-the-art (SOTA) efficient models such as DDistill-SR (Wang et al., 2022) or SAFMN (Sun et al., 2023) by a considerable margin, while utilizing only half or even less of the GMACS. Although our model is specifically designed for efficient SR, its scalability is evident as our larger model surpasses the SOTA lightweight transformer in performance while incurring lower computational costs. Overall, our key contributions are threefold:
  • We propose SeemoRe which matches the versatility of Transformer-based methods and the efficiency of CNN-based methods.
  • A Rank modulating expert (RME) is proposed to probe into the intricate inter-dependencies among the relevant feature projections in an efficient manner.
  • A Spatial modulating expert (SME) is proposed to integrate the complementary features extracted by SME by encoding the local contextual information.
2 Related Works
CNN-based SR.   In recent years, CNN-based techniques have outperformed traditional interpolation algorithms (Duchon, 1979) by learning a non-linear mapping between the input and target in an end-to-end training manner. The seminal SRCNN (Dong et al., 2014) introduced a three-layer convolutional approach for image super-resolution, later extended by works such as (Lim et al., 2017; Zhang et al., 2018b; Hui et al., 2019; Liang et al., 2021). VDSR (Kim et al., 2016) and EDSR (Lim et al., 2017) deepen networks using residual learning principles, with EDSR streamlining residual blocks for deeper training. Conversely, RCAN (Zhang et al., 2018a) introduces a novel residual-in-residual architecture for models exceeding 400 layers. While various spatial and channel attention mechanisms aim to enhance image reconstruction quality, CNN-based techniques still struggle to effectively utilize shared information across both dimensions. In this work, we aim to explore the interdependencies among the features in a computationally efficient way.
Transformer-based SR.   Thanks to its remarkable performance in high-level tasks (Dosovitskiy et al., 2021), the Transformer architecture has found its way into low-level vision tasks, such as image SR. Contemporary Transformer-based approaches aim to alleviate the computational load by confining self-attention to local regions and incorporating a higher degree of locality bias into their network design. SwinIR (Liang et al., 2021) incorporates local window self-attention and a shift mechanism inspired by the Swin Transformer design(Liu et al., 2021). Meanwhile, others like ELAN (Zhang et al., 2022) or ESRT  (Lu et al., 2022) reduce the feature dimensions by splitting or down-scaling to enhance the computational efficiency. Omni-SR (Wang et al., 2023) models pixel-interactions across different axes, creating universal correlations. SRFormer (Zhou et al., 2023) optimizes the computational efficiency by employing large window self-attention through the permutation of self-attention mechanisms. However, transformer-based methods typically demand significantly higher computational resources, even with smaller model capacities.
Figure 2: Architectute Overview. SeemoRe refines the feature representations via stacked Residual groups (RGs). Each RG consists of a Rank Modulating Exert (RME) and a Spatial Modulating Expert (SME). RME leverages the Mixeture of Low Rank (MoRE) to refine the global texture, while SME employs spatial enhancement experts (SEE) to supplement RME with the spatial cues.
Efficiency in SR.   In recent years, the pursuit of efficient SR techniques has gained significant momentum (Li et al., 2022; Ignatov et al., 2023; Li et al., 2023; Conde et al., 2023). Consequently, researchers have introduced streamlined neural architectures (Ignatov et al., 2021), network compression (Wang et al., 2022), reparameterization (Zhang et al., 2021b), and other training strategies to cater to the demand for efficiency. Initially, efficient SR methods utilized group convolutions and cascaded block designs to boost efficiency (Ahn et al., 2018; Hui et al., 2019). Subsequent advancements introduced convolution-based spatial or channel enhancement modules (Liu et al., 2020b). More recently, ShuffleMixer (Sun et al., 2022) integrates large kernel convolutions and feature shuffling, improving both computational efficiency and high-resolution reconstruction. SAFMN (Sun et al., 2023) improves the efficiency by collecting non-local features using a shallow pyramid. Despite improvements in several efficiency aspects brought up by the aforementioned approaches, there is still scope for a better trade-off between model efficiency and the restoration performance.
3 Methodology
In this section, we unveil the fundamental components of our proposed model tailored for efficient super-resolution. As demonstrated in Figure 2, our overall pipeline embodies a sequence of $N$ residual groups (RGs) and an upsampler layer. The initial step involves applying a 3$\times$3 convolution layer to generate the shallow features from the input low-resolution (LR) image. Subsequently, multiple stacked RGs are deployed to refine the deep features, easing the reconstruction of high-resolution (HR) images while maintaining efficiency. Each RG consists of a Rank modulating expert (RME) and a Spatial modulating expert (SME). Lastly, a global residual connection links the shallow features to the output of the deep features for capturing the high-frequency details and an up-sampler layer ( 3$\times$3 and pixel-shuffle (Shi et al., 2016)) is deployed for faster reconstruction.
3.1 Rank Modulating Expert
Unlike large kernel convolution (Hou et al., 2022) or self-attention (Vaswani et al., 2017) that rely upon resource-intensive matrix operations for modelling the LR-HR dependencies, we opt for modulating the most relevant interactions in low-rank in our quest for efficiency. Our proposed Rank modulating expert (RME) (see Figure 2) explores a Transformer alike architecture using Mixture of Low-Rank Expertise (MoRE) for modelling the relevant global informative features efficiently and a GatedFFN (Chen et al., 2023) for refined contextual feature aggregation.
Mixture of Low-Rank Expertise.   As illustrated in Figure 3, from a layer normalised input tensor $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$, we use a 3$\times$3 convolution for feature projection and then we split along the channel dimension to create two distinct views $\mathbf{x}_{a}~{}\text{and}~{}\mathbf{x}_{b}\in\mathbb{R}^{H\times W\times C}$. To efficiently aggregate the pixel-wise cross-channel context, we leverage a recursive strided convolution $t$ times followed by a refinement and upsampling step, resulting in the construction of the feature pyramid denoted as $\mathbf{\hat{x}}_{b}\in\mathbb{R}^{H\times W\times C}$. The process is formulated as follows:
$|p|_{\downarrow h\times w}$ $=\text{DConv}^{s}_{k\times k}(...(\text{DConv}^{s}_{k\times k}(\mathbf{x}_{b}))$ (1)
$\mathbf{\hat{x}}_{b}$ $=|~{}\mathbf{W}_{C\rightarrow C}(\text{DConv}_{3\times 3}(|p|_{\downarrow h\times w}))~{}|_{\uparrow H\times W},$ (2)
where $\text{DConv}_{k\times k}$ denotes a depth-wise convolution with kernel size $k$ and stride $s$, $\mathbf{W}_{C\rightarrow C}$ denotes a linear layer, $p$ represents the contextual feature pyramid. Simultaneously, a parallel depth-wise convolution extracts the local spatial context $\mathbf{\hat{x}}_{a}$ before feeding both the extracted feature maps into the mixture of low-rank expertise. This branched parallel design approach is chosen purposefully. In general, the downsampling of the feature maps impacts the reconstruction performance of SR methods. Therefore, we maintain the same resolution for general feature extraction while incorporating an additional path to capture global contextual cues efficiently, thereby circumventing any information loss.
To further delve into the intricacies of the inter-dependencies among the extracted features for reducing complexity, we deploy low-rank decomposition for the inputs while modeling the global contextual relationships. As demonstrated in Figure 3, a single low-rank expert ($\mathcal{E}$), takes as input the spatial features, $\mathbf{\hat{x}}_{a}$, and encoded pixel-wise contextual cues, $\mathbf{\hat{x}}_{b}$ and is formulated as:
$\mathcal{E}_{i}$ $=\mathbf{W}^{3}_{R_{i}\rightarrow C}(\mathbf{W}^{1}_{C\rightarrow R_{i}}\mathbf{\hat{x}}_{a}\odot\mathbf{W}^{2}_{C\rightarrow R_{i}}\mathbf{\hat{x}}_{b}),$ (3)
where the linear layers denoted as $\mathbf{W}_{C\rightarrow R_{i}}$, compress the encoded features along the channel dimension to their low-rank approximation $R_{i}$, where $i\in\{1,...,n\}$. After adeptly modulating the spatial cues through element-wise multiplication with the contextual cues in low-dimensional space, another linear layer $\mathbf{W}^{3}_{R_{i}\rightarrow C}$ extends the features back to the original dimension $C$ to extract the relevant channel-wise spatial content. Thereby, implicitly mixing the crucial spatial and channel dependencies in an efficient way.
However, manually determining the optimal low-rank ($R$) may not fully leverage all the inherent information for modulation, leading to underutilized model capacity. Thus, we employ a dynamic approach using a mixture of different low rank experts, with a routing network ($\mathcal{G}$) that systematically explores the search space to identify the ideal low-rank expert based on the input and network depth. Following (Shazeer et al., 2017), the final output $\mathbf{y}$ of the mixture of low-rank experts is as follows:
$\mathbf{y}$ $=\sum_{i}^{n}\mathcal{G}(\mathbf{\hat{x}}_{a})\mathcal{E}_{i}(\mathbf{\hat{x}}_{a},\mathbf{\hat{x}}_{b})+\mathbf{\hat{x}}_{a},$ (4)
where $\mathcal{G}(\cdot)$ and $\mathcal{E}_{i}(\cdot)$ denote the learned routing function and the output of the $i$-th expert, respectively. The sparsity inherent in the router function $\mathcal{G}(\cdot)$ optimizes computation by assigning greater weights to the top-$k$ low-rank experts. While at training time, our method learns from different experts, during inference, only the selected top-$k$ expert is utilized for computation, further enhancing the efficiency. More specifically, the inference complexity is not proportional to the number of experts. The design choices contributing towards the selection of the number of low-rank experts ($\mathcal{E}_{i}$) and the rank dimension ($R$) for memory efficient reconstruction is illustrated in Table 5 of the ablation study. We also provide the pseudocode for the proposed MoRE block in Algorithm 1.
Figure 3: Illustration of the proposed Mixture of Low-Rank Expertise (MoRE) as core block of the RME.
Algorithm 1 Mixture of Low-Rank Experise
1:  Input: Input feature $\mathbf{\hat{x}}_{a}$, semantic cues $\mathbf{\hat{x}}_{b}$
2:  Parameters: $n$ Experts $\mathcal{E}$, Router $\mathcal{G}$, Low-Rank dimensions $R_{i}=2^{i+1}$ with $i\in\{1,...,n\}$, top-k expert $k$
3:  Compute router outputs: $\mathbf{g}=\mathcal{G}(\mathbf{\hat{x}}_{a})$
4:  Normalize weights: $\mathbf{w}=\text{Softmax}(\mathbf{g})$
5:  Select top-k expert: $w_{k}=\text{topk}(\mathbf{w},k)$
6:  if training then
7:     for each $e\in\mathcal{E}$ do
8:        $\mathbf{y}^{i}_{e}=\mathbf{W}^{3}_{R_{i}\rightarrow C}(\mathbf{W}^{1}_{C\rightarrow R_{i}}\mathbf{\hat{x}}_{a}\odot\mathbf{W}^{2}_{C\rightarrow R_{i}}\mathbf{\hat{x}}_{b})$
9:     end for
10:     Compute final output: $\mathbf{y}=\sum_{i=1}^{n}w_{i}\cdot\mathbf{y}^{i}_{e}$
11:  else
12:     Compute final output: $\mathbf{y}=w_{k}\cdot\mathbf{y}^{k}_{e}$
13:  end if
14:  Output: Final output $\mathbf{y}$
3.2 Spatial Modulating Expert
We observe that the rank modulating expert is more dedicated towards investigating the global channel-wise contextual information, and its effectiveness would be complemented by the spatial-wise local information. Inspired by the previous work in classification (Yang et al., 2022; Hou et al., 2022), we design a spatial modulating expert (SME) (see Figure 2) comprising of a spatial enhancement expertise (SEE) block that efficiently captures the spatial-wise coupling followed by a GatedFFN (Chen et al., 2023) for feature refinement.
Spatial Enhancement Expertise.
While the vanilla self-attention (SA) mechanism (Vaswani et al., 2017) creates connections among all the input pixels, effectively capturing the relevant context, its quadratic computational complexity with image size poses limitations, particularly in high-resolution scenarios like image SR. Thus, our spatial enhancement expertise simplifies the computation of the similarity matrix $\mathbf{A}$ between keys $\mathbf{K}$ and queries $\mathbf{Q}$ by utilizing a striped depth-wise convolution with a large kernel, sequentially convolving the feature maps with $\mathbf{k_{1}}\in\mathbb{R}^{[1,k]}$ followed by $\mathbf{k_{2}}\in\mathbb{R}^{[k,1]}$. Specifically, we compute the locally enhanced spatial-wise features as follows:
$\mathbf{x}_{out}$ $=\text{DConv}^{s}_{k\times k}(\mathbf{W}^{4}_{C\rightarrow C}\mathbf{x}_{in})\odot\mathbf{W}^{5}_{C\rightarrow C}\mathbf{x}_{in},$ (5)
where $\odot$ is the Hadamard product, $\mathbf{W}^{4}_{C\rightarrow C}$ and $\mathbf{W}^{5}_{C\rightarrow C}$ are linear (project) layers, $\text{DConv}^{s}_{k\times k}$ denotes the striped depth-wise convolution, and $\mathbf{x}_{in}$ is the layer normalised output of the RME. The use of a large-kernel convolution facilitates a localized correlation among the pixels within the $k\times k$ window, emulating the window-based SA layers frequently employed in image restoration (Liu et al., 2021; Zamir et al., 2022; Chen et al., 2023), all the while preserving the efficiency benefits associated with convolutional layers as demonstrated in Table 4.
4 Experiments
Table 1: Comparison to efficient SR models. PSNR (dB normal-$\uparrow$) and SSIM (normal-$\uparrow$) metrics are reported on the Y-channel. Best and second best performances are highlighted. GMACS (G) are computed by upscaling to a $1280\times 720$ HR image. SeemoRe-T achieves state-of-the-art performance across all benchmarks with the lowest parameter count and computational demand. ‘-’ represents unreported results.
Method
Params GMACS Set5 Set14 BSD100 Urban100 Manga109
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
×$\times 2$
Bicubic
- - 33.66 .9299 30.24 .8688 29.56 .8431 26.88 .8403 30.80 .9339
CARN-M (Ahn et al., 2018)
412K 91 37.53 .9583 33.26 .9141 31.92 .8960 31.23 .9193 - -
IMDN (Hui et al., 2019)
694K 159 38.00 .9605 33.63 .9177 32.19 .8996 32.17 .9283 38.88 .9774
PAN (Zhao et al., 2020)
261K 71 38.00 .9605 33.59 .9181 32.18 .8997 32.01 .9273 38.70 .9773
DRSAN (Park et al., 2021)
370K 86 37.99 .9606 33.57 .9177 32.16 .8999 32.10 .9279 - -
DDistill-SR (Wang et al., 2022)
414K 128 38.03 .9606 33.61 .9182 32.19 .9000 32.18 .9286 38.94 .9777
ShuffleMixer (Sun et al., 2022)
394K 91 38.01 .9606 33.63 .9180 32.17 .8995 31.89 .9257 38.83 .9774
SAFMN (Sun et al., 2023)
228K 52 38.00 .9605 33.54 .9177 32.16 .8995 31.84 .9256 38.71 .9771
SeemoRe-T (ours)
220K 45 38.06 .9608 33.65 .9186 32.23 .9004 32.22 .9286 39.01 .9777
×$\times 3$
Bicubic
- - 30.39 .8682 27.55 .7742 27.21 .7385 24.46 .7349 26.95 .8556
CARN-M (Ahn et al., 2018)
415K 46 33.99 .9236 30.08 .8367 28.91 .8000 27.55 .8385 - -
IMDN (Hui et al., 2019)
703K 72 34.36 .9270 30.32 .8417 29.09 .8046 28.17 .8519 33.61 .9445
PAN (Zhao et al., 2020)
261K 39 34.40 .9271 30.36 .8423 29.11 .8050 28.11 .8511 33.61 .9448
DRSAN (Park et al., 2021)
410K 43 34.41 .9272 30.27 .8413 29.08 .8056 28.19 .8529 - -
DDistill-SR (Wang et al., 2022)
414K 57 34.37 .9275 30.34 .8420 29.11 .8053 28.19 .8528 33.69 .9451
ShuffleMixer (Sun et al., 2022)
415K 42 34.40 .9272 30.37 .8423 29.12 .8051 28.08 .8498 33.69 .9448
SAFMN (Sun et al., 2023)
233K 23 34.34 .9267 30.33 .8418 29.08 .8048 27.95 .8474 33.52 .9437
SeemoRe-T (ours)
225K 20 34.46 .9276 30.44 .8445 29.15 .8063 28.27 .8538 33.92 .9460
×$\times 4$
Bicubic
- - 28.42 .8104 26.00 .7027 25.96 .6675 23.14 .6577 24.89 .7866
CARN-M (Ahn et al., 2018)
415K 33 31.92 .8903 28.42 .7762 27.44 .7304 25.62 .7694 - -
IMDN (Hui et al., 2019)
715K 41 32.21 .8948 28.58 .7811 27.56 .7353 26.04 .7838 30.46 .9075
PAN (Zhao et al., 2020)
272K 28 32.13 .8948 28.61 .7822 27.59 .7363 26.11 .7854 30.51 .9095
DRSAN (Park et al., 2021)
410K 31 32.15 .8935 28.54 .7813 27.54 .7364 26.06 .7858 - -
DDistill-SR (Wang et al., 2022)
434K 33 32.23 .8960 28.62 .7823 27.58 .7365 26.20 .7891 30.48 .9090
ShuffleMixer (Sun et al., 2022)
411K 28 32.21 .8953 28.66 .7827 27.61 .7366 26.08 .7835 30.65 .9093
SAFMN (Sun et al., 2023)
240K 14 32.18 .8948 28.60 .7813 27.58 .7359 25.97 .7809 30.43 .9063
SeemoRe-T (ours)
232K 12 32.31 .8965 28.72 .7840 27.65 .7384 26.23 .7883 30.82 .9107
Table 2: Comparison to lightweight SR Transformers. SeemoRe-L outperforms or achieves comparable performance to compared Transformers while being more efficient. Best and second best performances are highlighted. ×$\times 3$ results are in the Supplemental.
Method
Params GMACS Set5 Set14 BSD100 Urban100 Manga109
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
×$\times 2$
Bicubic
- - 33.66 .9299 30.24 .8688 29.56 .8431 26.88 .8403 30.80 .9339
SwinIR-Light (Liang et al., 2021)
910K 244 38.14 .9611 33.86 .9206 32.31 .9012 32.76 .9340 39.12 .9783
ELAN-Light (Zhang et al., 2022)
621K 201 38.17 .9611 33.94 .9207 32.30 .9012 32.76 .9340 39.11 .9782
SRFormer-Light (Zhou et al., 2023)
853K 236 38.23 .9613 33.94 .9209 32.36 .9019 32.91 .9353 39.28 .9785
DAT-Light (Chen et al., 2023)
553K 194 38.24 .9614 34.01 .9214 32.34 .9019 32.89 .9346 39.49 .9788
SeemoRe-L (ours)
931K 197 38.27 .9616 34.01 .9210 32.35 .9018 32.87 .9344 39.49 .9790
×$\times 4$
Bicubic
- - 28.42 .8104 26.00 .7027 25.96 .6675 23.14 .6577 24.89 .7866
SwinIR-Light (Liang et al., 2021)
897K 64 32.44 .8976 28.77 .7858 27.69 .7406 26.47 .7980 30.91 .9151
ELAN-Light (Zhang et al., 2022)
601K 54 32.43 .8975 28.78 .7858 27.69 .7406 26.54 .7982 30.92 .9150
SRFormer-Light (Zhou et al., 2023)
873K 63 32.51 .8988 28.82 .7872 27.73 .7422 26.67 .8032 31.17 .9165
DAT-Light (Chen et al., 2023)
573K 50 32.57 .8991 28.87 .7879 27.74 .7428 26.64 .8033 31.37 .9178
SeemoRe-L (ours)
969K 50 32.51 .8990 28.92 .7888 27.78 .7428 26.79 .8046 31.48 .9181
Refer to caption Refer to caption
Urban100: img60 (×$\times 4$) Urban100: img73 (×$\times 4$) Urban100: img11 (×$\times 4$)
HR Crop
DDistill-SR
ShuffleMixer
SwinIR-Light
DAT-Light
SeemoRe-L
Figure 4: Visual comparison of SeemoRe with state-of-the-art methods on challenging cases for ×$\times 4$ SR from the Urban100 benchmark.
Table 3: Complexity Analysis. Runtime (ms, normal-$\downarrow$) and memory consumption (M, normal-$\downarrow$) averaged across 200 samples.
Method
Input Scale Time GPU Memory
DAT-Light
$[320,180]$$\times 4$ 210.12 8715.1
SwinIR-Light
131.25 6175.3
ShuffleMixer
7.40 1380.7
DDistill-SR
11.20 2822.1
SeemoRe-T (ours)
5.66 2744.9
[…]
Scale
Method
Params. GMACS Urban100 Manga109
×$\times 2$
[…]
191K 48 31.56 38.07
[…]
188K 49 31.60 38.59
SeemoRe-T 220K 45 32.22 39.01
[…]
Datasets and Evaluation.   Following the SR literature (Liang et al., 2021; Chen et al., 2023), we utilize DIV2K (Agustsson & Timofte, 2017) and Flickr2K (Lim et al., 2017) datasets for training. We produce LR images using bicubic downscaling of HR images. When testing our method, we assess its performance on canonical benchmark datasets for SR - Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010), BSD100 (Martin et al., 2001), Urban100 (Huang et al., 2015) and Manga109 (Matsui et al., 2017). We calculate PSNR and SSIM results on the Y-channel from the YCbCr color space.
Implementation Details.   We augment our training data with randomly extracted $64\times 64$-sized crops, with random rotation, horizontal and vertical flipping. Similar to (Sun et al., 2022, 2023), we minimize the L1-Norm between SR output and HR ground truth in the pixel and frequency domain using Adam (Kingma & Ba, 2017) optimizer for $500$K iterations with a batch size of $32$ and initial learning rate of $1\times 10^{-3}$ halving it at following milestones: [$250K$,$400K$,$450K$,$475K$]. All experiments are conducted with the PyTorch framework on NVIDIA RTX 4090 GPUs. We design our smallest model (SeemoRe-T) with 6 RGs. The feature dimension and channel expansion factor in GatedFFN are set to $36$ and $2$, respectively. For all MoRE sub-modules, we select an exponential growth of the channel dimensionality and choose in total of $3$ experts. The kernel size in SEE is set to $11\times 1$ 11. More details can be found in the supplemental, c.f. Table 6.
4.1 Comparison to State-of-the-Art Methods
We present quantitative results for ×$\times 2$, ×$\times 3$, and ×$\times 4$ image SR, comparing against current efficient state-of-the-art models in Table 1, including CARN-M (Ahn et al., 2018), IMDN (Hui et al., 2019), PAN (Zhao et al., 2020), DRSAN (Park et al., 2021), DDistill-SR (Wang et al., 2022), ShuffleMixer (Sun et al., 2022), and SAFMN (Sun et al., 2023). Additionally, we evaluate against lightweight variants of popular Transformer-based SR models such as SwinIR (Liu et al., 2021), ELAN (Zhang et al., 2022), and SRFormer (Zhou et al., 2023) in Table 2. Our proposed SeemoRe-T stands out as the most efficient method, consistently surpassing all other methods across all benchmarks and scale factors. For instance as clear from Table 1, on the Urban100 and Manga109 benchmarks (×$\times 2$), SeemoRe-T outperforms SAFMN (Sun et al., 2023) by $0.41$dB and $0.30$dB, respectively. Furthermore, with $47\%$ fewer parameters and $65\%$ fewer GMACS than DDistill-SR (Wang et al., 2022), SeemoRe-T achieves on average $0.12$ dB higher PSNR results across all benchmarks (×$\times 4$). Scaling our method up to a comparable size with lightweight Transformers yields comparable or superior results. As demonstrated in Table 2, our SeemoRe-L outperforms SwinIR-Light and SRFormer-Light on Manga109 (×$\times 4$) by $0.57$dB and $0.31$dB, while requiring fewer GMACS.
Visual Results.   We show visual comparisons (×$\times 4$) in Figure 4. In some challenging scenarios, the previous methods may suffer blurring artifacts, distortions, or inaccurate texture restoration. Contrary to others, our SeemoRe alleviates the blurring artifacts better and maintains structural fidelity. For instance, in image img60 and img73 from Urban100, certain methods like DDistill-SR, SwinIR-Light and DAT-Light fail to accurately reconstruct shadow patterns or window struts, whereas our method exhibits strong recovery of fine details. These visual comparisons highlight SeemoRe’s ability to reconstruct high-quality images by effectively leveraging local and contextual information. Coupled with quantitative comparisons, these findings underscore the effectiveness of our method. More visual results can be found in the Supplementary material.
4.2 Model Complexity Trade-Off
In the vision domain, scalability becomes more paramount. We strive to expand the limits of our SeemoRe framework, optimizing for both reconstruction fidelity and efficiency. The framework provides three complexity scales—tiny (T), base (B), and large (L)—with progressively improved reconstruction performance, c.f Figure 1. In Table 3, we present comparisons of memory usage and running time, demonstrating that our SeemoRe-T outperforms representative state-of-the-art methods. By using the low-rank feature modulation and simultaneous aggregation of the channel-spatial dependencies, the GPU consumption of our SeemoRe-T is $3\%$ less than DDistill-SR, while being $2$ times faster. Additionally, Table 3 highlights the significant efficiency advantage of SeemoRe over lightweight Transformers. Further results are provided in the Supplemental. To further underscore our method’s capability, we align SwinIR-Light and SRFormer-Light with a size and computational demand similar to ours, followed by retraining these downsized networks using our schedule. The results presented in Table 3 highlight that SeemoRe-T significantly outperforms both Transformer-based models by a considerable margin.
Table 4: Ablation on Blocks. GMACS (normal-$\downarrow$) are computed by upscaling to a $1280\times 720$ HR image.
Method
RME SME Params. GMACS Urban100 Manga109
Baseline
- - 232K 52 31.73 38.63
SeemoRe-T
$\checkmark$ - 249K 48 31.99 38.75
- $\checkmark$ 238K 53 32.05 38.96
$\checkmark$ $\checkmark$ 220K 45 32.22 39.01
[…]
Method
Kernel $k$ Params. GMACS Urban100 Manga109
SeemoRe-T
3 217K 44 32.04 38.84
5 218K 44 32.12 38.72
7 219K 45 32.10 38.96
11 220K 45 32.22 39.01
[…]
Table 5: Ablation on MoRE. Exponential growth yields best performance in terms of parameter counts and PSNR. $\#\mathcal{E}$ denotes the number of experts, and Dim. is the dimensionality of the rank.
Method $\#\mathcal{E}$ Dim. Params. Urban100 Manga109
SeemoRe-T
Growth
$i+2$
6 2,3,4,5,6,7 229K 32.12 39.01
$22*i+$ 2
4 2,4,6,8 224K 32.11 38.98
$2^{i}$
4 2,4,8,16 231K 32.21 39.02
$\#\mathcal{E}$
$2^{i}$
$3$ ${2,4,8}$ 220K 32.22 39.01
$2$ ${2,4}$ 214K 32.19 39.00
$1$ 22{2}2 211K 32.16 38.92
4.3 Ablation Study
We conduct detailed studies on the components within our approach. All experiments are conducted on the ×$\times 2$ setting.
Architecture Contribution.   As reported in Table 4, we evaluate the effectiveness of our proposed key architectural components by comparing them with a baseline model consisting solely of depthwise and pointwise convolutions, more details in Supplemental. After adding the proposed modules into the baseline model, their is a notable and persistent improvement in the results. The incorporation of RME or SME results in improvements of $0.26$ dB or $0.32$ dB on Urban100 over the baseline, respectively. Although both modules individually outperform the baseline with only a marginal increase in parameters, alternating the insertion of both the modules within each RG fully unleashes the model’s capabilities while enhancing the overall efficiency. Overall, our SeemoRe-T obtains a compelling gain of $0.49$ dB and $0.38$ dB on Urban100 and Manga109, respectively.
[…]
1234560050505050100100100100Residual Groups##\##Images×$\times 4$ Urban100123456Residual Groups×$\times 4$ Manga109
[…]
Figure 5: Low-Rank Analysis. (a) We plot the decisions made by the routing function for SeemoRe-T over the depth of the network.(b) We visualize the low-rank features of SeemoRe-T for ×$\times 4$ SR given example images from Urban100 and Manga109.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionInputBefore MoREAfter MoREAfter SEE
Figure 6: Feature Visualization. We present visualizations of feature maps before and after our proposed modules. Clearly, our MoRE block notably enhances activation sharpness via contextual feature modulation. Moreover, our SEE module improves learned representations by integrating spatial cues effectively.
Design choices of SME.   The main component of SME module, SEE deploys striped convolutions with large-kernel sizes to effectively module the spatial cues. Table 4 demonstrates that deploying large kernel sizes improves the overall performance of the model. In particular, the PSNR shows a notable gain of 0.18 dB on Urban100 dataset when increasing the kernel size from 3$\times$3 to 11$\times$11 (keeping other settings intact), with only 3K increase in parameters. It clearly proves that such a design benefits in the efficient use of the relevant information to augment the restoration of sharp regions spatially.
Design choices of RME.   We motivate our design choices for the MoRE module in RME by varying the growth function and the number of experts as depicted in Table 5. When pursuing a dynamic solution for determining the optimal low-rank dimensionality, it becomes necessary to design the corresponding search space. First, we present results for ×$\times 2$ upscaling on Urban100 and Manga109 using different growth functions. Based on the observed outcomes, it is evident that an exponentially increasing low-rank dimensionality yields the best performance with marginal increase in the parameters. Hence, we opt to retain this search space design in all further experimentation. Next, we analyze the reconstruction quality based on the number of experts in each MoRE module, while exponentially increasing the low-rank dimensionality. Based on these experiments, we assert that the efficient results are obtained when we have three, as our total number of experts.
5 Discussion on Experts
Our model integrates experts at varying levels, each specializing in crucial factors for SR. In this section, we aim to elucidate their expertise.
Mixture of Low-Rank Experts.   The decision-making process of the router at different network depths is illustrated in Figure 5. Notably, earlier blocks showcase a diverse range of rank choices ( $\mathcal{E}_{1}$,$\mathcal{E}_{2}$,$\mathcal{E}_{3}$), while deeper layers tend to favor lower ranks ($\mathcal{E}_{1}$) (Please note that for every $\mathcal{E}_{i}$, the corresponding rank dimension is $2^{i}$). This phenomenon can be attributed to the hierarchical feature learning nature of deep neural networks, aligning with our expectations. In fact, earlier layers typically capture low-level details and, at times, unwanted noise while reconstructing details in the input LR image, thus resulting in wide variations in the rank choices. In contrast, deeper layers focus on the main structures and key features required for SR. Hence, higher ranks at deeper layers are less favored, as they may introduce redundancy or noise that does not significantly contribute to the overall quality of the reconstructed image. This design aspect provides our method with the flexibility to adapt to the complexity of the task, a capability that, to the best of our knowledge, has not yet been explored in the image reconstruction community. In Figure 5, we further visualize the routing decisions and the corresponding low-rank feature maps for two exemplary input images. It is noteworthy that each individual rank carries distinct information while being mutually complementary. As the model depth increases, the network becomes proficient in restructuring these representations.
How important are MoRE and SEE?   To substantiate the significance of the proposed MoRE and SEE modules, we analyze the feature maps before and after integrating both blocks into the RME module as depicted in Figure 6. This analysis vividly showcases the advantages of leveraging MoRE for contextual information mining within RME. Notably, the activations exhibit reduced noise and enhanced sharpness. Additionally, we observe a synergistic interaction between MoRE and SEE at marked locations (indicated by red arrows): MoRE effectively refines global textures by filtering out noise, while SEE supplements over-filtered regions with critical local details.
6 Conclusion
We propose a novel ConvNet, named SeemoRe, for efficient and accurate image SR. Our SeemoRe excels in modeling local and contextual information, surpassing both previous CNN-based and lightweight Transformer approaches in terms of efficiency and reconstruction fidelity. Unlike other approaches, we empirically demonstrate both the scalability of efficiency and reconstruction performance. In our approach, we intricately design the rank modulation expert to discern the most pivotal features, enhancing this compressed representation with valuable contextual cues. Our spatial enhancement expert efficiently integrates local spatial-wise information, unlocking the full potential of our architecture. This novel approach optimally exploits the low information regime in the input image, enhancing detail reconstruction while improving efficiency. Extensive experiments on image SR demonstrate that our proposed SeemoRe achieves consistent superior performance over recent state-of-the-art efficient methods on all considered SR benchmarks, while even being on par with the lightweight Transformers in terms of reconstruction fidelity.
7 Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning, specifically efficient image super-resolution. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here. However, applying super-resolution methods in AI-assisted software raises ethical concerns about privacy invasion and increased surveillance capabilities. Adherence to transparency, accountability, and privacy rights is crucial to mitigate potential harm and ensure responsible deployment in alignment with societal values.
Appendix A Further Implementation Details
Table 6 outlines the architectural configurations and training settings employed to achieve the reported results in this study. Throughout all our experiments, we maintained a fixed random seed for reproducibility purposes. We based our implementation on the public PyTorch-based BasicSR [^1] https://github.com/XPixelGroup/BasicSR framework for architecture development and training. We use fvcore [^2] https://github.com/facebookresearch/fvcore Python package for computing GMACS and parameter counts.
Baseline for Architecture Contribution.
Here we provide more details about the baseline method for the ablation in Table 4. To ensure roughly equivalent parameter counts and computational complexity, we adopt $5$ RGs with a channel dimensionality of $48$. Within each RG, we integrate simple convolutional operators from our RME submodule without the MoRe module, while the SME module is simplified to a pointwise convolution.
Appendix B More Ablations
Optimization function.
In Table 8, we explore the impact of using the L1-Norm in FFT space to compare the model output with high-quality GT images. Compared to utilizing only the traditional L1 loss in RGB space, we observe an average performance improvement of $0.09$ dB on Urban100 and Manga109 datasets while using the combined losses. We acknowledge that only a few previous methods incorporate the same FFT loss (Sun et al., 2022, 2023); however, other efficient image super-resolution methods either employ a more intricate training schedule with multiple stages (Liu et al., 2020a; Kong et al., 2022) or utilize large-scale models for knowledge distillation (Wang et al., 2022).
Scaling the model size.
In Table 9, we detail the architecture, efficiency, and PSNR results across different model sizes on Urban100 and Manga109 datasets. Starting with SeemoRe-T, which has 220K parameters and 45 GMACS, each subsequent complexity stage doubles these figures. Notably, all model stages achieve state-of-the-art performance within their weight classes, with SeemoRe-L matching or even surpassing recent lightweight Transformer-based SR models.
Future work and limitations.
The proposed approach, employing a mixture of experts for feature modulation, is versatile for tasks with limited input information, such as low-light enhancement and denoising. Additionally, SeemoRe’s efficient design makes it a valuable solution for dynamic and resource-intensive environments. Expanding the number of experts in our network’s low-rank aspect poses challenges due to rapid feature dimensionality growth. Thus, our approach is currently limited to a small number of experts, contrasting with other fields leveraging larger expert ensembles.
Refer to caption Refer to caption Refer to caption
Manga109: img25 (×$\times 4$) Manga109: img99 (×$\times 4$) Manga109: img59 (×$\times 4$)
HR Crop
DDistill-SR
ShuffleMixer
SwinIR-Light
DAT-Light
SeemoRe-L
Figure 7: Visual comparison of SeemoRe with state-of-the-art methods on challenging cases for ×$\times 4$ SR from the Manga109 benchmark.
Refer to caption Refer to caption Refer to caption
Urban100: img04 (×$\times 4$) Urban100: img72 (×$\times 4$) Urban100: img92 (×$\times 4$)
HR Crop
DDistill-SR
ShuffleMixer
SwinIR-Light
DAT-Light
SeemoRe-L
Figure 8: Visual comparison of SeemoRe with state-of-the-art methods on challenging cases for ×$\times 4$ SR from the Urban100 benchmark.
Appendix C Visual Results.
We provide additional visual comparisons ($\times$4) in Figure 7 for the Manga109 benchmark and in Figure 8 for the Urban100 benchmark. Our SeemoRe framework consistently produces visually pleasing results, even on artistic images. In contrast to previous methods which exhibit flawed texture and character reconstruction, our proposed approach effectively reconstructs missing details, as illustrated in Figure 7, across all exemplary images considered. More concretely, when examining the example image img25 our SeemoRe network proficiently reconstructs the capital letter “I” within the text prompt “COMIC,” whereas SwinIR-Light and DAT-L encounter difficulty in producing any readable output. Additionally, in example image img04 our model significantly outperforms others in reconstructing the pattern with higher fidelity. Moreover, our model’s reconstruction of img92 in Figure 8 demonstrates reduced blurring and more distinct edges, enhancing overall visibility.